The data set below is on student absences and final grade. Sludem untudcat Student | Number of absences x | Final grade y (%)Assume the given below: x57 y=511, Σxy = 3745, Σ,--38993 2x2=579,

A. Find the linear correlation coefficient.

B. Find the least-squares regression line.

C.Using the model above, predict grade for 22 absences.

Note: Please, if you use any values from the tables, indicate which table they come from. Also, please write clearly on a piece of paper.
Answer #1

A.  Using the following formula to find out the correlation coefficient:

r =\frac{n.\sum xy - \sum x.\sum y}{\sqrt{\left [ n\sum x^{2}-(\sum x)^{2} \right ].\left [ n\sum y^{2}-(\sum y)^{2} \right ]}}

  7 3745-57511 v T7 * 579-57)21 · 17-389 93-511)21

.9442

B.  The independent variable is X, and the dependent variable is Y. The regression line is given by

y = a + b⋅x

In order to compute the regression coefficients, the following formula are used :

a= \frac{\sum y.\sum x^{2}-\sum x.\sum xy}{n.\sum x^{2}-(\sum x)^{2}}

  511 * 572-57 * 3745 7579 - 572

  102.49

b=

73745-57 511 7579- 572

ー-3.622

Substitute a and b in regression equation formula

y= 102.493 − 3.622⋅x

C. When, number of absence , x = 22

The grade y = 102.493 - 3.622*22

= 22.809